SUB-PICOSECOND INTERSYSTEM CROSSING AND VIBRATIONAL COOLING IN THE TRIPLET MAN-IFOLD OF 1-NITRONAPHTHALENE

\underline{CHRISTIAN REICHARDT}, R. AARON VOGT and CARLOS E. CRESPO-HERNÁNDEZ, Center for Chemical Dynamics, Department of Chemistry, Case Western Reserve University, 10900 Euclid Avenue, Cleveland, Ohio 44106.

The electronic energy relaxation of 1-nitronaphthalene (1NN) was studied in different solvents using broadband transient absorption spectroscopy with femtosecond time resolution. UV excitation of 1NN populates an unrelaxed S_{1}(ππ*) state, which decays by conformational relaxation (primarily twisting of the NO_{2} group) with a time constant of 100 fs. The twisting of the NO_{2} group and formation of a structurally relaxed singlet state opens up a doorway for ultrafast intersystem crossing (ISC) to a high-energy receiver triplet state T_{n}(nπ*), which then undergoes internal conversion to form a vibrationally excited T_{1}(ππ*) state. Quantum chemical calculations that include solvent effects support the experimental observations. Our results show that an essentially barrierless path connects the initial S_{1} state to the receiver T_{n} state, which enables the observation of vibrational energy transfer and its dependence on the surrounding solvent. According to this kinetic model, which was first proposed by Crespo-Hernández et al. for 1-nitropyrene[1], the S_{1}(ππ) electronic energy decays rapidly and irreversibly to dark triplet states, explaining why small nitro-polycyclic aromatic compounds are typically considered to be nonfluorescent.

Reference

[1] C. E. Crespo-Hernández, G. Burdzinski, R. Arce, J. Phys. Chem. A., 2008, 112,6313